3. Find the probability that a random sample of 12 senior high school students spends more than 24 hours. 


1
Expert's answer
2021-05-07T09:34:32-0400

Mean = 25

Standard deviation = 4

Z = (Mean - Score) / (Standard deviation / Square root of n)

Z = (24 - 25) / (4 / square root of 12)

Z = -0.87

Find probability using z distribution table.

P ( x > 24) = 1 - P (z > -0.87)

= 1 - 0.19

= 0.81 or 81%